OpenStudy (daisyblake11):

Solve, for θ in degrees 0° ≤ θ ≤ 360° : cotθ = 1

Nnesha (nnesha):

how would you write cot theta in terms of sin cos ?

OpenStudy (daisyblake11):

i really don't know

Nnesha (nnesha):

http://www.analyzemath.com/trigonometry/trigonometric_formulas.html bookmart this link http://prnt.sc/aj832u cot theta = ??

OpenStudy (daisyblake11):

i still dont get it

Nnesha (nnesha):

if you look at the pic cot theta = cos theta/sin theta \[\cot \theta =1 ~~ \rightarrow \frac{\cos \theta}{\sin \theta}=1\] now look at the unit circle at what degree cos divided by sin = 1 ?

OpenStudy (daisyblake11):

360 ?

OpenStudy (daisyblake11):

0 actually

Nnesha (nnesha):

x-coordinate = cos and y-coordiante represent sin so at 0 degree cos =1 and sin =0 so when you divide cos /sin 1/0 isn't equal to 1 1/0 is undefined

Nnesha (nnesha):

http://i.stack.imgur.com/r8uHr.gif

OpenStudy (daisyblake11):

45 degrees

Nnesha (nnesha):

yep that's one what else ? :)

OpenStudy (daisyblake11):

225 degrees

Nnesha (nnesha):

yep

OpenStudy (daisyblake11):

so the answers are 45 and 225 degrees ?

Nnesha (nnesha):

yep

OpenStudy (daisyblake11):

you are amazing. thank you so much

Nnesha (nnesha):

haha thanks:) yw